# Minimalist Market Design: A Framework for Economists with Policy Aspirations

### v3

@@ 11.4 Teaming Up with Three Leaders in Bioethics
The system recommended in the Original Guidance was basically a simplified version of the priority point system of White et al. (2009), with only two ethical principles instead of three. Naturally, we wondered how the Boston Globe could possibly make such a major error in its coverage. Subsequently, we observed that parts of the Original Guidance used ambiguous language, mentioning possible preference given to medical personnel who are vital to treating others, even though the final recommendation did not reflect this consideration.131131131The ambiguous language in the Original Guidance was later removed in a Revised Guidance issued on April 20, 2020, possibly due to an inquiry by our team about this inconsistency to Robert Truog, who was a member of the working group mentioned in the Boston Globe story.
@@
Although our solution was not adopted for the crisis allocation of ventilators, over the next year, it was implemented for other scarce medical resources such as antiviral therapies, monoclonal antibodies, and vaccines, thanks to our collaboration with our new partners.
@@ 11.4.1 Policy Impact on Allocation of COVID-19 Therapies in Pennsylvania
Under the leadership of Douglas White, the UPMC system was designed by our team as a special case of a reserve system and was implemented in Western Pennsylvania during the shortage period of Remdesivir (White et al., 2022). The software for the system was also developed and provided by our core team of three design economists.
@@ 11.5.2 NASEM Framework for Equitable Vaccine Allocation
Richmond and Henrico Reserve System for Phase 1b COVID-19 Vaccine Rollout
Categories, their share of COVID-19 vaccines, and factors that affect category-specific priorities were specified by the health districts as follows:132132132Category-specific priorities are only partially specified in RichmondaCity (2021).
@@
    People ages 16-64 with Comorbidities (23% of units)
@@ 11.6 Policy Impact on Allocation of mAb Therapies in Massachusetts
Massachusetts Reserve System for Allocation of mAb Therapeutics
Categories, their respective share of mAb infusions, factors influencing category-specific priorities, and the processing sequence of categories were specified by the state as follows:
@@ 12 Liver Exchange
As discussed in detail earlier in Section 7.7, the core idea in IKE revolves around mitigating welfare losses that occur when kidney patients directly receive living donations from blood-type-compatible donors who are not blood-type-identical (e.g., a blood-type A patient with a blood-type O donor). By participating in kidney exchange instead, these patients can help additional individuals receive transplants while still securing one for themselves. IKE introduces an “institutional” incentive for such patients to participate in kidney exchange by granting them priority on the deceased-donor (DD) list should a future failure of their transplant kidney occur. Its non-minimalist nature stems from the need to integrate two separate “institutions,” namely, the allocation mechanisms for DD kidneys and living-donor kidneys, which are typically managed independently in the U.S. and elsewhere.
However, as the welfare gains are explicitly tied to an “institutional” incentive mechanism within IKE, realizing them poses greater challenges. If only we could discover a “biological” incentive mechanism inherent to the transplantation procedure, perhaps we would have better luck transforming this idea into practice.
In the mid-2010s, we finally discovered one such “biological” incentive mechanism we had been seeking—albeit for another organ: the liver. Little did I know that not only would this discovery pave the way for a series of subsequent breakthroughs, but, in an initiative that became uniquely close to my heart, it would also materialize into a program—the Banu Bedestenci Sönmez Liver Paired Exchange System—dedicated to my beloved wife, whom we tragically lost not long after our discovery (BBS-LPE, 2024).
@@ 12.1 Living-Donor Liver Transplantation and Liver Paired Exchange
In principle, donor exchange can be implemented for any organ where living-donor transplantation is possible, and the liver is one such organ. Despite potential donors having only one liver (unlike the kidney), it is often possible to donate a portion of it—typically either the right lobe, the left lobe, or a part of the left lobe known as Segment 2-3.
@@
In the early 2000s, living-donor kidney transplants in the U.S. surpassed 6,000 annually. In contrast, after reaching a peak of 524 in 2001, the number of living-donor liver transplants (LDLTs) sharply declined in the subsequent years (Ratnera and Sandoval, 2010):
@@
The successful implementation of liver exchange in these two centers served as compelling proof of concept, prompting us to pay closer attention to this potential application of market design. However, in the early 2010s, the annual number of LDLTs consistently decreased in the U.S., reaching a minimum of 205 in 2012. Indeed, not long after the liver exchanges from South Korea and Hong Kong were reported, another donor death made headlines in the U.S. (Cohen, 2012).
Even though we were not fully convinced that the conditions were ripe to pursue a project in liver exchange, these developments still compelled us to learn about the risks of partial hepatectomy, the surgery in which part of the liver is removed from a donor. It was then that we made a key discovery—a “biological” incentive mechanism that could motivate compatible pairs to participate in liver exchange. Building on this idea, as well as another concept driven by a liver transplant technique called “dual-lobe liver transplant,” we extensively researched liver exchange throughout much of the mid-2010s as part of a team that included Haluk Ergin, Utku Ünver, and me (Ergin et al., 2017, 2020).
@@ 12.2 A Biological Incentive Mechanism for Liver Exchange
To minimize the risk of death during or immediately after a liver transplant operation, an adult patient must receive a liver graft comprising at least 40% of their dysfunctional liver’s size. Receiving a graft smaller than this limit significantly elevates the risk of patient death due to the small-for-size syndrome. Given that the liver’s weight is approximately 2% of the body weight, meeting this criterion necessitates a graft-to-recipient weight ratio (GRWR) of at least 0.8%. In principle, eligible donors can contribute either their larger right lobe—typically 60-75% of their liver—or the left lobe—typically 25-40% of their liver. However, these figures imply that the left lobe of a donor is often too small for a patient. Consequently, in most cases, eligible donors can only donate their right lobes to their adult patients.
@@
Another crucial statistic reveals that, in much of the world, donor mortality and morbidity rates are much higher for right-lobe hepatectomy than for left-lobe hepatectomy (Lee, 2010; Mishra et al., 2018).133133133Lee (2010) reports mortality rates of 0.4-0.5% for right hepatectomy and 0.1% for left hepatectomy. Mishra et al. (2018) reports morbidity rates of 28% for right hepatectomy and 7.5% for left hepatectomy. Consequently, while the left lobe is frequently deemed unsuitable for transplantation, it represents a much safer option for the donor. This observation forms the basis for a “biological” incentive that may compel blood-type-compatible patient-donor pairs to participate in liver exchange.
@@
Consider a scenario involving a blood type A patient with a petite blood type O donor. The patient can only receive the larger right lobe of the donor. In a typical liver exchange pool, there are often numerous blood type O patients who are incompatible with their corresponding blood type A donors.134134134In matching theory, we refer such pairs as “under-demanded” (Roth et al., 2005b, 2007). Frequently, some of them—pediatric or slender individuals—require smaller liver grafts. While the first patient is compatible with their donor, unlike the case of a kidney transplant, a donor exchange between the two pairs becomes a visibly appealing option. This is due to the fact that the first donor can undergo a much less risky left-lobe hepatectomy through liver exchange, as opposed to the considerably higher-risk direct donation to their own patient.
@@
Hence, the requirement for size compatibility in liver transplantation, combined with the varying risks associated with right-lobe and left-lobe hepatectomy, establishes a compelling “biological” incentive for numerous compatible pairs. Expanding upon this insight and concentrating on logistically simpler 2-way exchanges, our work in Ergin et al. (2020) formulated and proposed an efficient and incentive-compatible liver exchange mechanism. The decision to focus on 2-way exchanges was compelling as, until 2022, no center had undertaken the logistically more challenging 3-way or larger liver exchanges.
@@
By utilizing this mechanism and leveraging aggregate statistics from South Korea, we demonstrated not only the potential to increase the number of LDLTs by over 30% but also the ability to raise the proportion of less risky left-lobe transplants. All that remained was to find a center willing to pursue this idea. In 2019, we identified such a center in our homeland, Turkey.
@@ 12.3 Living-Donor Liver Transplantation at Malatya, Turkey
The transplant team and facilities at İnönü University grew so significantly that, in 2018, they performed three LDLTs simultaneously. Despite their substantial accomplishments and the extensive experience accumulated by the team, they remained relatively unknown throughout much of Turkey, let alone the world, as they were located in a less developed part of the country.
In 2019, a distinguished surgeon and the director of a prominent Ivy League university hospital in the U.S. visited the Malatya team. After learning about their 2018 experience with three simultaneous LDLTs, he suggested that İnönü University promote this remarkable feat by applying for a Guinness World Record. Once this idea resonated with university leadership, Yılmaz shared that, by then, his team had developed the capacity to perform five simultaneous liver transplants.
@@
However, not everyone shared the enthusiasm for achieving a world record through five simultaneous liver transplants. Following formal complaints submitted to the Health Ministry by a world-renowned Turkish transplant surgeon—the founding father of transplantation in Turkey—seeking disciplinary action against the group, a number of reporters also heavily criticized the Malatya team. While government officials remained supportive of the team, realizing that their bid for a Guinness World Record was misguided, Yılmaz withdrew the application for the record.
@@
To my disappointment, no one in the Malatya group responded to my second message for two months. Unaware of the developments regarding the Guinness World Record application and the subsequent publicity crisis that unfolded a month after my initial correspondence with the group, I decided in early July 2019 to follow up on my earlier email while visiting my parents in Turkey. I informed Yılmaz that I was in the country and would be delighted to visit his team in Malatya if an in-person meeting would be more informative. Yılmaz promptly responded to my message, explaining that he hadn’t been able to get back to me earlier due to his heavy schedule and providing his phone number. He expressed that they would be very happy to host me in Malatya. I immediately called him and arranged a visit with their team a few days later.
@@ 12.4 How a Personal Tragedy Influenced My Policy Efforts in Turkey
My initiative in Malatya wasn’t my first policy effort in Turkey. Ever since my failed attempt to reform the centralized Turkish college admissions mechanism in 1997 (see Section 6.6), I have made several other attempts. Ünver and I made quite a few of our initial discoveries in kidney exchange while working at Koç University in the early 2000s, fueling our enthusiasm to see our innovations save lives in our homeland.
The aspiration to save lives through these ideas was a primary motivation for our decision to leave Koç University in 2004 and move to the U.S. This significant life choice was made possible by the considerable sacrifice of my wife, Banu, whom I had married only two years earlier. Prior to our relocation, Banu was a successful dentist in İstanbul and had never contemplated leaving Turkey. By choosing to move abroad, she became a homemaker in a foreign country, leaving behind her career and social life to support my aspirations to improve the world through my research.
@@
During this period, I conducted extensive research on breast cancer and, drawing on insights from Banu’s decade-long treatments, I even attempted to initiate a project on cancer clinical trial design. Unfortunately, this project did not materialize. While Banu spent her last few months with us in hospice at home between April and August 2016, partly as a coping mechanism, I occasionally dedicated time to our liver exchange project. The potential of this project to save lives was the sole motivation for me to engage in any research during such a challenging time. In that period, the most arduous in my life, an unexpected development unfolded.
@@
For about a year, we interacted with bureaucrats at the Turkish Health Ministry to arrange various details for the planned system. By the end of June 2017, the final form of a protocol between the Health Ministry and our team of two design economists was ready. However, just as we prepared to sign the protocol, Akdağ’s tenure as Health Minister expired in July 2017. We never heard from the Health Ministry bureaucrats after this date, and the protocol was never signed.
@@
12.5 Forging a Partnership with a Renowned Liver Transplant Team
@@
At the same time, considering the remarkable professional and logistical capacity of the İnönü University liver transplant team to conduct five simultaneous transplants, and the publicity crisis they faced while trying to highlight this capacity, I felt that this time could be very different.
@@
In Malatya, I encountered a very friendly group eager to learn about the liver exchange system we envisioned. After a productive discussion with their leader, Sezai Yılmaz, I delivered a comprehensive presentation to the entire team. During my visit, I introduced the (never signed) protocol drafted by the Health Ministry two years earlier to establish a national system in honor of my beloved wife, Banu. Yılmaz proposed that they could readily adapt this protocol for a single-center system and promptly arranged a meeting with Prof. Dr. Ahmet Kızılay, the President of İnönü University. With enthusiastic support from the president, we reached an agreement to establish and jointly manage the Banu Bedestenci Sönmez Liver Paired Exchange (BBS-LPE) System.

### v4

@@ 11.4 Teaming Up with Three Leaders in Bioethics
The system recommended in the Original Guidance was basically a simplified version of the priority point system of White et al. (2009), with only two ethical principles instead of three. Naturally, we wondered how the Boston Globe could possibly make such a major error in its coverage. Subsequently, we observed that parts of the Original Guidance used ambiguous language, mentioning possible preference given to medical personnel who are vital to treating others, even though the final recommendation did not reflect this consideration.145145145The ambiguous language in the Original Guidance was later removed in a Revised Guidance issued on April 20, 2020, possibly due to an inquiry by our team about this inconsistency to Robert Truog, who was a member of the working group mentioned in the Boston Globe story.
@@
Although our solution was not adopted for the crisis allocation of ventilators, over the next year it was implemented for other scarce medical resources—such as antiviral therapies, monoclonal antibodies, and vaccines—thanks to our collaboration with our new partners (see Table 6 for notable reserve systems adopted during the COVID-19 pandemic).
Table 6: Notable reserve systems adopted during the COVID-19 pandemic.
  • Note: Reproduced from Pathak et al. (2024), Management Science, licensed under $C$ BY 4.0.
@@ 11.4.1 Policy Impact on Allocation of COVID-19 Therapies in Pennsylvania
Under the leadership of Douglas White, the UPMC system was designed by our team as a special case of a reserve system and was implemented in Western Pennsylvania during the shortage period of Remdesivir (White et al., 2022). The software for the system was also developed and provided by our core team of three economic designers.
@@ 11.5.2 NASEM Framework for Equitable Vaccine Allocation
Richmond and Henrico Reserve System for COVID-19 Vaccine Rollout.
In their Phase 1b allocation, the Richmond and Henrico health districts gave the following specification of categories, their respective shares of COVID-19 vaccines, and the factors shaping category-specific priorities:
@@
    People ages 16–64 with Comorbidities (23% of units)
@@ 11.6 Policy Impact on Allocation of mAb Therapies in Massachusetts
Massachusetts Reserve System for Allocation of mAb Therapeutics.
In the Massachusetts Reserve System for the Allocation of mAb Therapeutics, the state specified the categories, their shares of infusions, the factors influencing category-specific priorities, and the processing sequence of categories:
@@ 12 Liver Exchange
As discussed in detail earlier in Section 7.7, the core idea of IKE centers on mitigating welfare losses that occur when kidney patients receive living donations from blood-type-compatible—but not blood-type-identical—donors (for example, a blood-type A patient with a blood-type O donor). By participating in kidney exchange instead, these patients can help others receive transplants while still securing one for themselves. IKE introduces an “institutional” incentive for such patients to participate in kidney exchange by granting them priority on the deceased-donor (DD) list should a future failure of their transplant kidney occur. Its non-minimalist nature stems from the need to integrate two separate “institutions,” namely, the allocation mechanisms for DD kidneys and living-donor kidneys, which are typically managed independently in the U.S. and elsewhere.
However, as the welfare gains from IKE are explicitly tied to an “institutional” incentive mechanism, realizing them poses greater challenges. If only we could discover a “biological” incentive mechanism inherent to the transplantation procedure, perhaps we would have better luck transforming this idea into practice.
In the mid-2010s, we finally discovered one such “biological” incentive mechanism we had been seeking—though for another organ: the liver. Little did I know that this discovery would not only pave the way for a series of subsequent breakthroughs but also, in an initiative uniquely close to my heart, culminate in a program—the Banu Bedestenci Sönmez Liver Paired Exchange System—dedicated to my beloved wife, whom we tragically lost not long after our discovery (BBS-LPE, 2025).
@@ 12.1 Living-Donor Liver Transplantation and Liver Paired Exchange
In principle, donor exchange can be implemented for any organ where living-donor transplantation is possible, and the liver is one such organ. Despite potential donors having only one liver (unlike the kidney), it is often possible to donate a portion of it—typically either the right lobe, the left lobe, or a part of the left lobe known as Segments 2-3 (see Figure 31).
Figure 31: Liver anatomy. Segments 1-4 form the left lobe and Segments 5-8 form the right lobe. In some cases, Segments 2–3 of the left lobe can be transplanted into pediatric patients. Abbreviation: IVC: inferior vena cava; PV: portal vein. Reproduced from Orcutt et al. (2016), Frontiers in Surgery, licensed under $C$ BY 4.0.
@@
In the early 2000s, living-donor kidney transplants in the U.S. surpassed 6,000 annually. By contrast, after peaking at 524 in 2001, the number of living-donor liver transplants (LDLTs) declined sharply in the following years (Ratnera and Sandoval, 2010).146146146After falling to an all-time low of 219 cases in 2009, the annual number of LDLTs gradually rose over the next 25 years, reaching a record high of 658 in 2023 (OPTN, 2025).
@@
The successful implementation of liver exchange in these two centers served as compelling proof of concept, encouraging us to pay closer attention to this potential application of market design. Yet, in the early 2010s, the annual number of LDLTs in the U.S. remained consistently in the 200s. Indeed, not long after the liver exchanges from South Korea and Hong Kong were reported, the death of another donor made headlines in the U.S. (Cohen, 2012).
Even though we were not fully convinced that the conditions were ripe to pursue a project in liver exchange, these developments still compelled us to learn about the risks of partial hepatectomy, the surgery in which part of the liver is removed from a donor. It was then that we made a key discovery—-a “biological” incentive mechanism that could, at least in theory, motivate compatible pairs to participate in liver exchange. Building on this idea, as well as another concept driven by a liver transplant technique called “dual-lobe liver transplant,” we extensively researched liver exchange throughout much of the mid-2010s as part of a team that included Haluk Ergin, Utku Ünver, and me (Ergin et al., 2017, 2020).
@@ 12.2 A Biological Incentive Mechanism for Liver Exchange
To minimize the risk of death during or immediately after liver transplantation, an adult patient must receive a liver graft comprising at least 40% of the size of their dysfunctional liver. Receiving a graft smaller than this limit significantly elevates the risk of patient death due to the small-for-size syndrome. Given that the liver’s weight is approximately 2% of body weight, meeting this criterion necessitates a graft-to-recipient weight ratio (GRWR) of at least 0.8%. In principle, eligible donors can contribute either their larger right lobe (typically 60–75% of their liver) or the left lobe (typically 25–40% of their liver) (see Figure 31). However, these constraints imply that the left lobe of a donor is often too small for an adult patient. Consequently, in most cases, eligible donors can only donate their right lobe to adult patients.147147147For donor safety, the remnant liver after hepatectomy needs to be at least 30% of the original liver volume. This means right lobes larger than 70% of the liver cannot be safely transplanted.
@@
Another crucial statistic reveals that, in much of the world, donor mortality and morbidity rates are much higher for right-lobe hepatectomy than for left-lobe hepatectomy (Lee, 2010; Mishra et al., 2018).148148148Lee (2010) reports mortality rates of 0.4–0.5% for right hepatectomy and 0.1% for left hepatectomy. Mishra et al. (2018) reports morbidity rates of 28% for right hepatectomy and 7.5% for left hepatectomy. Consequently, while the left lobe is frequently deemed unsuitable for transplantation, it represents a much safer option for the donor. This observation forms the basis for a “biological” incentive that may compel blood-type-compatible patient–donor pairs to participate in liver exchange.
@@
Consider a scenario involving a blood type A patient with a petite blood type O donor. The patient can only receive the larger right lobe of the donor. In a typical liver exchange pool, there are often numerous blood type O patients who are incompatible with their corresponding blood type A donors.149149149In matching theory, we refer to such pairs as “under-demanded” (Roth et al., 2005b, 2007). Frequently, some of these patients—pediatric or slender individuals—require smaller liver grafts. While the first patient is compatible with their donor, unlike in kidney transplantation, a donor exchange between the two pairs becomes a visibly appealing option. This is because the first donor can undergo a much less risky left-lobe hepatectomy through liver exchange, as opposed to the considerably higher-risk direct donation to their own patient.
@@
Thus, the need for size compatibility in liver transplantation, together with the differing risks of right- versus left-lobe hepatectomy, creates a strong “biological” incentive for many compatible pairs. Building on this insight, our work in Ergin et al. (2020) developed an efficient and incentive-compatible liver exchange mechanism centered on 2-way exchanges—the only type attempted before 2022, since no center had yet undertaken the more complex 3-way or larger exchanges.
@@
By utilizing this mechanism and leveraging aggregate statistics from South Korea, even with only 2-way exchanges, we demonstrated not only the potential to increase the number of LDLTs by over 30% but also the ability to raise the proportion of less risky left-lobe transplants. All that remained was to find a center willing to pursue this idea. In 2019, we identified such a center in our homeland, Turkey.
@@ 12.3 Living-Donor Liver Transplantation at Malatya, Turkey
The transplantation team and facilities at İnönü University grew so significantly that, in 2018, they performed three LDLTs simultaneously. Despite their substantial accomplishments and the extensive experience accumulated by the team, they remained relatively unknown throughout much of Turkey, let alone the world.
In 2019, a distinguished surgeon and the director of a prominent Ivy League university hospital in the U.S. visited the Malatya team. After learning about their 2018 experience with three simultaneous LDLTs, he suggested that İnönü University promote this remarkable feat by applying for a Guinness World Record. Once this idea resonated with the university leadership, Yılmaz informed him that, by then, his Malatya team had developed the capacity to perform five simultaneous liver transplants.
@@
However, not everyone shared the enthusiasm for achieving a world record through five simultaneous liver transplants. Following formal complaints submitted to the Health Ministry by a world-renowned Turkish transplant surgeon—the founding father of transplantation in Turkey—seeking disciplinary action against the group, a number of reporters also heavily criticized the Malatya team. While government officials remained supportive of the team, Yılmaz recognized that the Guinness World Record attempt risked overshadowing the accomplishment itself and withdrew the application.
@@
To my disappointment, no one in the Malatya group replied to my second message for two months. Unaware of the developments surrounding the Guinness World Record application and the publicity crisis that followed a month after my initial correspondence, I decided in early July 2019—while visiting my parents in Turkey—to follow up on my earlier email. I wrote to Yılmaz that I was in the country and would be delighted to visit his team in Malatya if an in-person meeting would be more helpful. He responded promptly, apologizing for the delay due to his heavy schedule, sharing his phone number, and expressing that they would be very happy to host me. I called him right away and arranged a visit with his team a few days later.
@@ 12.4 How a Personal Tragedy Influenced My Policy Efforts in Turkey
My initiative in Malatya wasn’t my first policy effort in Turkey. Ever since my failed attempt to reform the centralized Turkish college admissions mechanism in 1997 (see Section 6.6), I have made several other attempts. Ünver and I made quite a few of our initial discoveries in kidney exchange while working at Koç University in the early 2000s, fueling our enthusiasm to advance these innovations in our homeland.
The aspiration to save lives through these ideas worldwide was a primary motivation for our decision to leave Koç University in 2004 and move to the U.S. This significant life choice was made possible by the considerable sacrifice of my wife, Banu, whom I had married only two years earlier. Prior to our relocation, Banu was a successful dentist in İstanbul and had never contemplated leaving Turkey. By choosing to move abroad, she transformed my professional aspirations into our joint dream, and became a homemaker in a foreign country, leaving behind her career and social life to support my efforts to improve the world through research.
@@
During this period, I conducted extensive research on breast cancer and, drawing on insights from Banu’s decade-long treatments, I even attempted to initiate a project on cancer clinical trial design. Unfortunately, this project did not materialize. While Banu spent her last few months with us in hospice at home between April and August 2016, partly as a coping mechanism, I occasionally dedicated time to our liver exchange project. The potential of this project to save lives was the sole motivation for me to engage in any research activity during such a challenging time. In that period, the most arduous in my life, an unexpected development unfolded.
@@
For about a year, we interacted with bureaucrats at the Turkish Health Ministry to arrange various details for the planned system. By the end of June 2017, the final form of a protocol between the Health Ministry and our team of two economic designers was ready. However, just as we prepared to sign the protocol, Akdağ’s tenure as Health Minister expired in July 2017. We never heard from the Health Ministry bureaucrats after this date, and the protocol was never signed.
@@
12.5 Forging a Partnership with a Renowned Liver Transplantation Team
@@
At the same time, considering the remarkable professional and logistical capacity of the İnönü University liver transplantation team to conduct five simultaneous transplants, and the publicity crisis they faced while trying to highlight this capacity, I felt that this time could be very different.
@@
In Malatya, I encountered a very friendly group eager to learn about the liver exchange system we envisioned. After a productive discussion with their leader, Sezai Yılmaz, I delivered a comprehensive presentation to the entire team. During my visit, I shared with them the (never signed) protocol drafted by the Health Ministry two years earlier to establish a national system in honor of my beloved wife, Banu. Yılmaz proposed that they could readily adapt this protocol for a single-center system and promptly arranged a meeting with Prof. Dr. Ahmet Kızılay, the President of İnönü University. With enthusiastic support from the president, we reached an agreement to establish and jointly manage the Banu Bedestenci Sönmez Liver Paired Exchange (BBS–LPE) System.